OpenStudy (anonymous):

Write 5.6 as a percent. 56% 5.6% 560% 0.56%

OpenStudy (princeharryyy):

kindly check your options again @kawaii-chan

OpenStudy (misty1212):

multiply it by 100

OpenStudy (misty1212):

\[5.6\times 100=?\]

OpenStudy (anonymous):

ok thanks misty1212

OpenStudy (asnaseer):

remember that percentage means "out of 100", so if we want 5.6 as a percentage we need to find x such that:\[5.6=\frac{x}{100}\]

OpenStudy (aripotta):

writing 5.6 into a percent would be 5.6/100 which should be 0.056%, but that's not one of your choices

OpenStudy (princeharryyy):

@AriPotta :p

OpenStudy (anonymous):

OK result is 5.6*100 = 560 %

OpenStudy (mathstudent55):

@AriPotta You interpreted the question as "how do you write 5.6% as a decimal?", but your answer is incorrect. 5.6% as a decimal is 5.6/100 = 0.056 (no percent sign). The question the original poster asked is: "how do you write 5.6 as a percent?" The answer is 5.6 * 100% = 560%

OpenStudy (aripotta):

aahh i see. i misread the question. i'm sorry!